OpenStudy (anonymous):

The number of cars at a dealership dropped from 64 to 48 after a weekend sale. What is the percent of decrease in the number of cars? A. 20% B. 25% C. 33% D. 75%

OpenStudy (31356):

No. of cars dropped from 64 to 48 No. of cars before= 64 No cars after weekend sale= 48 Percentage decrease 64−4864×100=1664×100 =161644×100=25%

OpenStudy (31356):

one percent of 64 = 0.64 one houndred percent of 64 = 64 so 100 times 0.64 = 64 then for 48 cars we want percentage times 0.64 = 48 or percentage = 48 divided by 0.64 then 48 is 75% of 64 and the decrease in percentages is 25

OpenStudy (31356):

Thanks for the medal!

OpenStudy (anonymous):

thank you :) can you help me with another one ?

OpenStudy (anonymous):

The population in a small city increased from 1250 people to 2780 people. What is the percent of increase in the number of people? Round to the nearest percent. A. 45% B. 55% C. 122% D. 153%

OpenStudy (31356):

122.39999999999999% round it up to 122.4%

OpenStudy (31356):

There you go.

OpenStudy (anonymous):

The number of dogs in a pet store decreased from 35 to 31. What is the percent of decrease in the number of dogs? Round to the nearest percent. A. 11% B. 13% C. 40% D. 97%

OpenStudy (anonymous):

I would probably go with 13.

OpenStudy (anonymous):

Are you in connections?